OpenStudy (anonymous):

Please select two choices below to earn full credit. Which two numbers below are equal? [Select two of the numbers below.] A. 1 3/15 B. 1 1/5 C. 3 1/15 D. 13/15

OpenStudy (alexandervonhumboldt2):

A B

OpenStudy (alexandervonhumboldt2):

3/15=1/5 1=1 so A and B

OpenStudy (alexandervonhumboldt2):

@mjperez18

OpenStudy (anonymous):

Okay thank you!